OpenStudy (anonymous):

Can someone explain this problem to me? tan(-20°) = -tan20° True- The answer is true, but I dont know why. False

OpenStudy (harsimran_hs4):

ok let`s try can you tell sin(-x) = ? (in terms of sin(x))

OpenStudy (anonymous):

umm Im not sure

OpenStudy (harsimran_hs4):

anyways try...tell whatever is in your mind

OpenStudy (anonymous):

sin(180 -) = sin

OpenStudy (harsimran_hs4):

ok so sin(-x) = - sin(x) (clear ?)

OpenStudy (anonymous):

okay

OpenStudy (harsimran_hs4):

now try for cos(-x) = ?

OpenStudy (anonymous):

okay cos(-x)=-cos(x)

OpenStudy (harsimran_hs4):

NO!!

OpenStudy (anonymous):

cos(-x)=cos(x)

OpenStudy (harsimran_hs4):

so now on right track so tan(-x) = sin(-x)/ cos(-x) = -sin(x)/ cos(x) = - tan(x) do you get it now?

OpenStudy (anonymous):

yes, I believe so

OpenStudy (harsimran_hs4):

why are you still not sure what`s the confusion?

OpenStudy (anonymous):

Im saying I believe so, because Im new to this so im not going to know this, but I believe I know enough to get through this lesson

OpenStudy (harsimran_hs4):

ok so does that end this question or there is still something that you need to ask?

OpenStudy (anonymous):

No Im good, Thank you for your time!

OpenStudy (harsimran_hs4):

:)